HUBBLE SPACE TELESCOPE DAILY REPORT #1778 PERIOD COVERED: 0000Z (UTC) 12/12/96 - 0000Z (UTC) 12/13/96 Daily Status Report as of 348/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6348 (Blue Continuum Snapshots of 3CR Radio = Galaxies) The WF/PC-2 was used to make some snapshot images of the blue = continuum in radio galaxy 3C411.0. Combined with existing data, these = exposures will provide high spatial resolution information on the color = distribution within galaxies, redding due to dust, spacial indices of the = optical synchrotron jets, and colors of the active nuclei and tidal jets. = The observations were executed as scheduled, and no anomalies were = noted. 1.2 Completed FOS/BL/GHRS/1 6918 (FOS Cycle 6: Internal/External = Offset) The Faint Object Spectrograph (Blue Side) and the High Resolution = Spectrograph (Side One) were both used to make observations of standard = star HD207757 in order to make internal and external calibration = measurements. Offsets between internal and external wavelength scales = were measured for several different gratings. The observations were = executed as scheduled, and no anomalies were noted. 1.3 Completed FOS/RD/GHRS/1 6911 (FOS Cycle 6: Discriminator Test) The Faint Object Spectrograph (Red Side) and the High Resolution = Spectrograph (Side One) were used to make internal flat and a dark = calibration images as part of a Cycle 6 optimal discriminator setting = test. This must be done because the noise and gain are both known to be = temperature sensitive. Observations were also taken of standard star = HD207757. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed GHRS/1 6575 (Observations of Interstellar Clouds in the = Galactic Halo) The High Resolution Spectrograph (Side One) was used to make = ultraviolet spectrographic observations of the halo star HD119608 in = order to study the foreground interstellar medium absorption lines. = These spectra will be used to determine the physical properties of the = individual clouds, and will provide more information on the elemental = abundances, the relation between the ionized and neutral gas, and the = distribution of the various species. The observations were completed as = planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: Aside from several temperature limit violations which normally = occur at this time of the year, there were no reported operational = problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH